OpenStudy (anonymous):

how can i integrate ∫x*(sin(x))^2

OpenStudy (anonymous):

Intergration by parts

OpenStudy (anonymous):

let x be u and sinx be dv

OpenStudy (anonymous):

formula is uv-int(v.du)

OpenStudy (jamesj):

Wait a second. I think the question is \[ \int x \sin^2 x \ dx \] yes?

OpenStudy (lgbasallote):

wouldnt assigning u to sin^2(x) be easier? it's a little confusing integrating sin^2(x)...im just suggesting i dunno which is easier in the end..jsyk

OpenStudy (anonymous):

can you give me all serial solution

OpenStudy (jamesj):

The first thing to is to get rid of the sin^2 term. Use a trig identity with cos(2x).

OpenStudy (anonymous):

standard trick is to write what jamesj said

OpenStudy (jamesj):

cos(2x) = cos^2 x - sin^2 x = 1 - 2.sin^2 x Thus sin^2 x = ....

OpenStudy (anonymous):

i couldnot solve it

OpenStudy (jamesj):

\[ \sin^2 x = \frac{1}{2} (1 - \cos 2x) \] Hence \[ \int x . \sin^2 x \ dx = \frac{1}{2} \int (x - x.\cos 2x) \ dx \] See what to do now?

OpenStudy (anonymous):

go on

OpenStudy (jamesj):

So you don't know what to do next? Surely you know how how to integrate the first part, \[ \int x \ dx \] yes? Now, for the second part, use integration by parts.

OpenStudy (anonymous):

thanks